# Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems

### v2

@@ Abstract
A quantum kinetic theory of the linear response to an electric field is provided from a controlled expansion of the Keldysh theory at leading order, for a multiband electron system with weak scalar disorder. The response is uniquely partitioned into intraband distribution functions and interband quantum coherences. A new explicit formula is provided for the latter, outlining their true nature as local slave quantities, for which the mesoscopic gradients of the former is uncovered as a source term, opening a new research area. The precise connection with the Kubo formula with vertex corrections in the ladder approximation is established. A pedagogical application to spin-orbit torque theory in the two dimensional electron gas demonstrates the striking efficiency of the new formalism. Some important implications for the theory of orbital transport are also briefly discussed.
@@
In this Letter, we provide a quantum kinetic theory of the linear response to a nearly constant and uniform electric field ${\bm{E}}$, for a multiband electron system, in presence of weak scalar disorder. With close connections with similar derivations in high-energy physics [41, & references therein], this is obtained from a controlled semiclassical expansion of the Schwinger-Keldysh theory [42, 43, 44], combined with a gauge invariant quantum geometric decomposition [45]. With the details the derivation to be found here [46], we expose our theory in terms of intraband and interband responses, uncovering mesoscopic gradients of the former as a previously overlooked source of the latter, which opens new areas of research. In the homogeneous infinite case, using Matsubara formalism [15, 16] , we establish the equivalence of our results with the Kubo formula with vertex correction in the ladder approximation. As a pedagogical example, we derive a theory of the spin-orbit torque (SOT) in the two-dimensional electron gas (2DEG), and uncover a new instance of intrinsic response as a Fermi surface property, reminiscent of the classical result obtained by Haldane in the case of the AHE [47]. Finally, we outline some implications for the theory of the recently observed orbital edge accumulation [48, 49] and orbital torque [50, & references therein], and we conclude by mentioning some future extensions.
@@
while the intraband part reads $\hat{\rho}_{intra}^{\scriptscriptstyle(1)}=\sum_{n}f^{\scriptscriptstyle(1)}_{n}\hat{P}_{n}$  [^1] This form of $\hat{\rho}^{\scriptscriptstyle(1)}$ is strictly valid for non degenerate bands. In the case of extended degeneracy, which is most generically realized in the form of Kramer’s degeneracy in time reversal and inversion symmetric systems, it remains valid if spin-orbit interaction can be safely neglected. Otherwise, a traceless part shall also be considered. In addition, in Eq.(1) we omit principal value terms which are usually interpreted as disorder-induced renormalizations of the Hamiltonian. Both can be found in [46]. and is controlled by a system of linear BEs
@@
$W_{nq}\equiv W_{nq}({\bm{p}},\underaccent{\tilde}{\bm p})=\frac{2\pi n_{i}v_{0}^{2}}{\hbar}{\rm Tr}\left[\hat{P}_{n}\underaccent{\tilde}{\hat{P}}_{q}\right].$ (4)
@@
It has to be noted that we systematically use a tilde under-accent to indicate that a certain quantity depends on the integration momentum variable $\underaccent{\tilde}{\bm p}$. After introducing $\tau\propto\hbar/(n_{i}v_{0}^{2})$ as the relaxation time scale, we see that $\xi=1/\tau$ defines a natural scaling parameter for the disorder strength, while $\hbar$ is of course the formal expansion parameter for the quantum corrections to the classical kinetics [41, 45, & references therein].
@@
While the commonly considered decomposition of the Kubo-Streda type formulas into Fermi sea and Fermi surface terms [5, 7] is not physically significant in itself [47], the quantum kinetic approach leads naturally to a unique and gauge invariant split at the Wigner function level i.e., $\hat{\rho}^{\scriptscriptstyle(1)}=\hat{\rho}^{\scriptscriptstyle(1)}_{intra}+\hat{\rho}^{\scriptscriptstyle(1)}_{inter}$. Its physical meaning is further illuminated by Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems), establishing the interband response components as slave variables locally determined by the quantum geometry and the solutions of the BE (2). Hence, the quantum kinetic formulation establishes that interband contributions to observable densities are not quantities susceptible to advective or diffusive transport in real space. Considering the different terms on the right-hand side (RHS) of Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems), we recognize the first term as of purely quantum geometric origin and of order $\hbar\xi^{0}$, which will yield for instance the intrinsic contribution to the AHE [58, 11] which can be expressed in terms of the Berry curvature. If we look now at the third term on the RHS of our Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems), since $f_{n}^{\scriptscriptstyle(1)}\sim 1/\xi$, as it can be immediately verified by direct inspection of Eqs.(2-4), we see that it is also of order $\hbar\xi^{0}$ due to the compensating factor $(n_{i}v_{0}^{2})/\hbar\propto\xi$ in front of the momentum integral. Hence, even if this contribution is disorder mediated, it is ultimately independent of its strength. As demonstrated in the next section, this corresponds to the vertex correction in the ladder approximation to the interband part of the Kubo formula, for which the quantum kinetic approach provides an explicit expression at the Wigner function level, in terms of the solutions to the BE (2). As previously mentioned, even recent theoretical studies rely most often on a simplified version of the Kubo-Streda formula at the CBA level, with the leading disorder correction to the interband response found to be of order $\hbar\xi$ [59], which is qualitatively wrong and totally misses the true leading effect of disorder. The second term on the RHS of Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems) indicates that mesoscopic gradients in spatially inhomogeneous solutions of the BE (2) are also a source of interband quantum coherences, for which we provide here an explicit expression. To the best of our knowledge, this is an entirely new finding. Such gradients typically appear near boundaries and interfaces on the length scale of the electron mean-free-path $l_{e}\propto 1/\xi$ [60, 61, 62]. This opens a new area of research, since one can envision to engineer such gradients using nanofabrication techniques. Since $\hbar\partial_{\bm{x}}f_{n}^{\scriptscriptstyle(1)}\sim(\hbar/l_{e})f_{n}^{\scriptscriptstyle(1)}\sim\hbar\xi^{0}$, this is of the same order as the other contributions. Regarding the intraband response, it shall also be noted that there is no Berry curvature corrections to the BE (2), at the leading order $f_{n}^{\scriptscriptstyle(1)}\sim 1/\xi$. There are however sub-leading corrections at order $\hbar\xi^{0}$ which involve the quantum geometry, and are related to the intrinsic side-jump and skew-scattering contributions found with the semiclassical approach [35, 36, 37]. This is beyond the scope of the present letter, since a complete account requires to go beyond the ladder approximation [63], and will be provided within our general framework in a future work.
@@
Pedagogical application to the SOT in the 2DEG– As a pedagogical example, we address now the microscopic theory of the SOT in the weakly disordered 2DEG, a topic previously investigated with the Kubo formula [20, 65]. The model Hamiltonian is
@@
with ${\bm{p}}\equiv(p^{x},p^{y})$, $\varepsilon_{0}$ some paramagnetic band dispersion, $J_{e}>0$ the exchange constant, ${\bm{m}}$ a classical unit vector aligned with the magnetization and ${\bm{b}}_{\scriptscriptstyle SO}$ some momentum dependent SO field with $b_{\scriptscriptstyle SO}\ll J_{e}$. The operator $\hat{\bm{\sigma}}$ is the vector of Pauli matrices, while $\hat{\sigma}_{0}$ is the identity matrix in spin space. The Hamiltonian (11) is a generalization, to arbitrary band and SO field dispersions, of the model considered in [20]. If we neglect internal dissipation [66, 67], the dynamics of ${\bm{m}}$ is controlled by the Landau-Lifschitz equation [68, 69] i.e., $(M_{\scriptscriptstyle S}/\gamma)\partial_{t}{\bm{m}}={\bm{m}}\times(\delta\mathcal{F}/\delta{\bm{m}})\equiv{\bm{T}}$, with $\gamma(<0)$ the gyromagnetic ratio, $M_{\scriptscriptstyle S}$ the saturation magnetization, $\mathcal{F}$ the thermodynamic potential and ${\bm{T}}$ the torque density. The SOT, in linear response to a quasi uniform and constant electric field, is then given by ${\bm{T}}^{(1)}={\bm{m}}\times(\delta\mathcal{F}^{\scriptscriptstyle(1)}/\delta{\bm{m}})$, with $\mathcal{F}^{\scriptscriptstyle(1)}$ the linear change in $\mathcal{F}$. By application of the Hellmann-Feynman theorem [70, 71] at finite temperature [72], and assuming a spatially uniform magnetizartion, we then obtain ${\bm{T}}^{\scriptscriptstyle(1)}={\bm{m}}\times\langle\delta\hat{h}/\delta{\bm{m}}\rangle^{\scriptscriptstyle(1)}=J_{e}{\bm{m}}\times\langle\hat{\bm{\sigma}}\rangle^{\scriptscriptstyle(1)}$. The band energies and associated projectors for the Hamiltonian (11) are trivially obtained as $\varepsilon_{\sigma}=\varepsilon^{0}+\sigma\Delta$ and $\hat{P}_{\sigma}=\frac{1}{2}\left(\hat{\sigma}_{0}+\sigma{\bm{u}}\cdot\hat{\bm{\sigma}}\right)$, respectively ; in which we have introduced $\sigma\in\{+1,-1\}\equiv\{\uparrow,\downarrow\}$, with $\Delta\approx J_{e}(1+{\bm{m}}\cdot\underaccent{\bar}{\bm b}_{\scriptscriptstyle SO})$ and ${\bm{u}}\approx{\bm{m}}+{\bm{m}}\times(\underaccent{\bar}{\bm b}_{\scriptscriptstyle SO}\times{\bm{m}})$, as we are approximating from now on all quantities at leading order in $\underaccent{\bar}{\bm b}_{\scriptscriptstyle SO}={\bm{b}}_{\scriptscriptstyle SO}/J_{e}$. The interband Berry connection reads $\hat{\mathcal{A}}^{i}_{\sigma\underaccent{\tilde}{\sigma}}\ \approx\ \frac{{\bf i}\underaccent{\tilde}{\sigma}}{4}\partial_{p^{i}}\underaccent{\bar}{\bm b}_{\perp}\cdot\hat{\bm{\sigma}}\ +\ \frac{1}{4}({\bm{m}}\times\partial_{p^{i}}\underaccent{\bar}{\bm b}_{\perp})\cdot\hat{\bm{\sigma}}$, with $\underaccent{\bar}{\bm b}_{\perp}={\bm{m}}\times(\underaccent{\bar}{\bm b}_{\scriptscriptstyle SO}\times{\bm{m}})$. The intrinsic contribution to the SOT, from the first term on the RHS of Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems), immediately ensues as
@@
${\bm{T}}^{\scriptscriptstyle(1)}_{int}\approx\frac{\hbar}{2}\int_{p}\left(f^{\scriptscriptstyle(eq)}_{\uparrow}-f^{\scriptscriptstyle(eq)}_{\downarrow}\right)(e{\bm{E}})\cdot\partial_{\bm{p}}\underaccent{\bar}{\bm b}_{\perp}.$ (12)
@@
After recognizing that ${\bm{E}}\cdot\partial_{\bm{p}}\underaccent{\bar}{b}_{\perp}^{i}=\partial_{\bm{p}}\cdot(\underaccent{\bar}{b}_{\perp}^{i}{\bm{E}})$ and assuming the system to be degenerate, an application of the divergence theorem to the RHS of Eq.(12) yields
@@
with $dl_{p}$ being the infinitesimal line element along the 1D Fermi surface for band $\sigma$ in the absence of SO interaction i.e., $FS_{\sigma}$ 222 If the Fermi surface intersects the Brillouin zone (BZ) boundary, the line integrals over some portions of this boundary resulting from the application of the divergence theorem does not contribute to the RHS of Eq.(12). Indeed, since the 2D BZ is by definition a primitive unit cell of the reciprocal space over which $\hat{h}$ is periodic, such contributions will cancel out by pairs, from pairs of line segments on opposite sides of the BZ boundary, one being the translation of the other by a primitive reciprocal lattice vector. Hence, we have established that under the stated hypothesis, the intrinsic SOT, which according to Eq.(12) would be classified as a Fermi sea contribution, can equally be expressed as a Fermi surface one according to Eq.(13). This exactly parallels the classical result obtained by Haldane in the case of the non quantized part of the AHE conductivity [47]. Having established this non trivial result, we focus now on the disorder correction to ${\bm{T}}^{\scriptscriptstyle(1)}_{int}$, which is stemming from the third term on the RHS of our Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems). Specializing it to the case at hand, we readily obtain this contribution to the interband Wigner function as
@@
Implications for the theory of orbital effects – Regarding the theory of out-of-equilibrium orbital angular momentum (OAM) generation in time reversal and inversion symmetric metals, as an explanation for several recently observed effects [50, 48, 49], most Kubo formula based studies have focused so far on the computation of the intrinsic OHE conductivity in the bulk [14, & references therein], irrespective of the chosen definition for the OAM observable. In our formulation, the intrinsic OHE current can be recovered from the first term on the RHS of Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems). Beyond the possible fragility of this current under disorder corrections [21], which can be straightforwardly assessed from the third term on the RHS of our Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems), we want to point out again the importance of the second term i.e., of quantum coherences generated by the gradients. It can be shown [74] that this is a source of orbital edge accumulation, that is only partially correlated to the OHE conductivity through the Berry connection and which may be an important explanatory factor to the previously cited experimental observations.
@@
Conclusions – We provide a quantum kinetic theory of the linear response to an electric field slowly varying in time and space, for multiband electron systems in presence of weak scalar disorder. We uncovered gradients in the intraband response on mesoscopic scales as a source of interband coherence, opening a new research area as it suggests alternate ways to locally modulate current induced responses e.g., spin or orbital accumulation, by controlling the boundaries shape or scattering properties. While our theory constitutes an alternative to the Kubo formula in the stated context, with distinct advantages for finite systems and in going beyond the CBA, we have proven the equivalence of the two formulations at the ladder approximation level in the homogeneous case. We applied our approach to SOT theory for a simple model of 2DEG, recovering and generalizing with striking efficiency a major vertex correction, while uncovering a new instance of intrinsic response as a Fermi surface property. We also suggest possible implications for the theory of orbital effects. Our formalism is applicable to realistic material models, up to ab-initio generated tight-binding Hamiltonian, through numerical resolution of the Boltzmann equation. We are already undertaking such studies in relation to SOT in van der Waals magnets with $C_{3}v$ symmetry [75], a subject of high current interest [18]. Finally, the disorder model can be easily extended, for instance to include SO scattering contributions, and we also intend to address sub-leading contributions to the intraband response in a future work.

### v3

@@ Abstract
A quantum kinetic theory of the linear response to an electric field is provided from a controlled expansion of the Keldysh theory at leading order, for a multiband electron system with weak scalar disorder. The response is uniquely partitioned into intraband distribution functions and interband quantum coherences. A new explicit formula is provided for the latter, outlining their true nature as local dependent quantities, for which the mesoscopic gradients of the former is uncovered as a source term, opening a new research area. The precise connection with the Kubo formula in the ladder approximation is established. A pedagogical application to spin-orbit torque theory in the two dimensional electron gas demonstrates the striking efficiency of the quantum kinetic approach. Some important implications for the theory of orbital transport are also briefly discussed.
@@
In this Letter, we provide a quantum kinetic theory of the linear response to a nearly constant and uniform electric field ${\bm{E}}$, for a multiband system, in presence of weak scalar disorder. We combine high-energy physics techniques [41, & references therein], with a gauge invariant formulation of quantum geometry [42], leading to a much streamlined, more rigorous, way of performing the semiclassical expansion of the Keldysh theory [43, 44, 45], as detailed in the Supplemental Material [46] (see also references [47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67] therein). Our theory uniquely partitions the response into intraband and interband contributions, uncovering mesoscopic gradients of the former as a previously overlooked source of the latter. In the uniform case, using Matsubara formalism [15, 16], we establish the equivalence of our results with the Kubo formula in the ladder approximation. An application to spin-orbit torque (SOT) theory in the two-dimensional electron gas (2DEG) demonstrates the distinct advantages of our approach, and uncover a new instance of intrinsic response as a Fermi surface property, reminiscent of the classical result obtained by Haldane in the case of the AHE [68]. Finally, we outline some implications for the theory of the recently observed orbital edge accumulation [69, 70] and orbital torque [71, & references therein], and we conclude by mentioning some future extensions.
@@
while the intraband part reads $\hat{\rho}_{intra}^{\scriptscriptstyle(1)}=\sum_{n}f^{\scriptscriptstyle(1)}_{n}\hat{P}_{n}$  [^1] This form of $\hat{\rho}^{\scriptscriptstyle(1)}$ is strictly valid for non degenerate bands. In the case of extended degeneracy, which is most generically realized in the form of Kramer’s degeneracy in time reversal and inversion symmetric systems, it remains valid if the spin-orbit interaction can be safely neglected. Otherwise, a traceless matrix part shall also be considered. In addition, in Eq.(1) we omit principal value terms which are usually interpreted as disorder-induced renormalization terms of the Hamiltonian. Both can be found in [46]. and is controlled by a system of linear BEs
@@
$W_{nq}\equiv W_{nq}({\bm{p}},\underaccent{\tilde}{\bm p})=\frac{2\pi n_{i}v_{0}^{2}}{\hbar w_{n}}{\rm Tr}\left[\hat{P}_{n}\underaccent{\tilde}{\hat{P}}_{q}\right],$ (4)
@@
with $w_{n}\geq 1$ being the degree of degeneracy of band $n$, and with a tilde under-accent indicating throughout this Letter that a certain quantity depends on momentum $\underaccent{\tilde}{\bm p}$. After introducing $\tau\propto\hbar/(n_{i}v_{0}^{2})$ as the relaxation time scale, we see that $\xi=1/\tau$ defines a natural scaling parameter for the disorder strength, while $\hbar$ is of course the formal expansion parameter for the quantum corrections to the classical kinetics [41, 78, & references therein].
@@
While the commonly considered decomposition of the Kubo-Streda type formulas into Fermi sea and Fermi surface terms [5, 7] is not physically significant in itself [68], the quantum kinetic approach leads naturally to a unique and gauge invariant split at the Wigner function level i.e., $\hat{\rho}^{\scriptscriptstyle(1)}=\hat{\rho}^{\scriptscriptstyle(1)}_{intra}+\hat{\rho}^{\scriptscriptstyle(1)}_{inter}$. Its physical meaning is further illuminated by Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems), establishing the interband response components as dependent quantities locally determined by the quantum geometry and the solutions of the BE (2). Hence, the quantum kinetic formulation provides the key new physical insight that interband contributions to observable densities are not quantities susceptible to advective or diffusive transport in real space. Looking at the right-hand side (RHS) of Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems), we recognize the first term as of purely quantum geometric origin and of order $\hbar\xi^{0}$, which yields for instance the intrinsic contribution to the AHE [79, 11] controlled by the Berry curvature in reciprocal space. If we look now at the third term on the RHS of our Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems), since $f_{n}^{\scriptscriptstyle(1)}\sim 1/\xi$, as it can be immediately verified by direct inspection of Eqs.(2-4), we see that it is also of order $\hbar\xi^{0}$ due to the compensating factor $(n_{i}v_{0}^{2})/\hbar\propto\xi$ in front of the momentum integral. Hence, even if this contribution is disorder mediated, it is ultimately independent of its strength. As demonstrated in the next section, this corresponds to the vertex correction in the ladder approximation to the interband part of the Kubo formula, for which we provide here an explicit and compact expression, a substantial advantage compared to the Kubo formula approach for concrete calculations. Moreover, as previously mentioned, even recent works rely most often on a simplified version of the Kubo-Streda formula at the CBA level, with the leading disorder correction to the interband response found to be of order $\hbar\xi$ [80], which is qualitatively wrong and totally misses the true leading effect of disorder. The second term on the RHS of Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems) indicates that mesoscopic gradients in spatially inhomogeneous solutions of the BE (2) are also a source of interband quantum coherences, for which we provide here an explicit expression. To the best of our knowledge, this is an entirely new finding. Such gradients typically appear near boundaries and interfaces on the length scale of the electron mean-free-path $l_{e}\propto 1/\xi$ [81, 82, 83]. This opens a new area of research, since one can envision to engineer such gradients using nanofabrication techniques. Since $\hbar\partial_{\bm{x}}f_{n}^{\scriptscriptstyle(1)}\sim(\hbar/l_{e})f_{n}^{\scriptscriptstyle(1)}\sim\hbar\xi^{0}$, this is of the same order as the other contributions. Regarding the intraband response, it shall also be noted that there is no Berry curvature corrections to the BE (2), at the leading order $f_{n}^{\scriptscriptstyle(1)}\sim 1/\xi$. There are however sub-leading terms at order $\hbar\xi^{0}$ i.e., in particular anomalous velocity, intrinsic side-jump and skew-scattering, previously found with the semiclassical approach [35, 36, 37]. This is beyond the scope of the present letter, since a complete account requires to go beyond the ladder approximation [84], and will be provided within the same general framework in a future work.
@@
Pedagogical application to the SOT in the 2DEG– As a pedagogical example, and as a way to demonstrate the advantages of our kinetic formulation when performing concrete calculations beyond the CBA, we address now the microscopic theory of the SOT for a weakly disordered 2DEG in presence of exchange splitting induced by some classical magnetization vector, a topic previously investigated with the Kubo formula [20, 86]. The model Hamiltonian is
@@
with ${\bm{p}}\equiv(p^{x},p^{y})$, $\varepsilon_{0}$ some paramagnetic band dispersion, $J_{e}>0$ the exchange constant, ${\bm{m}}$ a classical unit vector aligned with the magnetization and ${\bm{b}}_{\scriptscriptstyle SO}$ some momentum dependent SO field with $b_{\scriptscriptstyle SO}\ll J_{e}$. The operator $\hat{\bm{\sigma}}$ is the vector of Pauli matrices and corresponds here to the electron spin operator, while $\hat{\sigma}_{0}$ is the identity matrix in spin space. The Hamiltonian (11) is a generalization, to arbitrary band and SO field dispersions, of the model considered in [20]. If we neglect dissipation [87, 88], the dynamics of ${\bm{m}}$ is controlled by the Landau-Lifschitz equation [89, 90] i.e., $(M_{\scriptscriptstyle S}/\gamma)\partial_{t}{\bm{m}}={\bm{m}}\times(\delta\mathcal{F}/\delta{\bm{m}})\equiv{\bm{T}}$, with $\gamma(<0)$ the gyromagnetic ratio, $M_{\scriptscriptstyle S}$ the saturation magnetization, $\mathcal{F}$ the thermodynamic potential and ${\bm{T}}$ the torque density. The SOT, in linear response to a quasi uniform and constant electric field, is then given by ${\bm{T}}^{(1)}={\bm{m}}\times(\delta\mathcal{F}^{\scriptscriptstyle(1)}/\delta{\bm{m}})$, with $\mathcal{F}^{\scriptscriptstyle(1)}$ the linear change in $\mathcal{F}$. By application of the Hellmann-Feynman theorem [91, 92] at finite temperature [93], and assuming a spatially uniform magnetization, we then obtain ${\bm{T}}^{\scriptscriptstyle(1)}={\bm{m}}\times\langle\delta\hat{h}/\delta{\bm{m}}\rangle^{\scriptscriptstyle(1)}=J_{e}{\bm{m}}\times\langle\hat{\bm{\sigma}}\rangle^{\scriptscriptstyle(1)}$. Hence we see that the SOT is controlled by the component of the expectation value of the spin density perpendicular to the classical magnetization, in linear response to the field. In the case of an SO field breaking inversion symmetry, one expects both intraband and interband contributions to the SOT [94]. Here, we uniquely focus on the interband contribution, for which a previous Kubo calculation has established in some special cases the cancellation of the intrinsic SOT by the vertex correction [20].
We start by deriving anew the intrinsic SOT i.e., the contribution of purely quantum geometric origin. The band energies and associated projectors for the Hamiltonian (11) are trivially obtained as $\varepsilon_{\sigma}=\varepsilon^{0}+\sigma\Delta$ and $\hat{P}_{\sigma}=\frac{1}{2}\left(\hat{\sigma}_{0}+\sigma{\bm{u}}\cdot\hat{\bm{\sigma}}\right)$, respectively ; in which we have introduced $\sigma\in\{+1,-1\}\equiv\{\uparrow,\downarrow\}$, with $\Delta\approx J_{e}(1+{\bm{m}}\cdot\underaccent{\bar}{\bm b}_{\scriptscriptstyle SO})$ and ${\bm{u}}\approx{\bm{m}}+{\bm{m}}\times(\underaccent{\bar}{\bm b}_{\scriptscriptstyle SO}\times{\bm{m}})$, in which the almost equal symbol is used from now on to indicate quantities approximated at leading order in the SO field normalized by the exchange energy i.e., $\underaccent{\bar}{\bm b}_{\scriptscriptstyle SO}={\bm{b}}_{\scriptscriptstyle SO}/J_{e}$, which is a natural small parameter of the problem. The interband Berry connection reads $\hat{\mathcal{A}}^{i}_{\sigma\underaccent{\tilde}{\sigma}}\ \approx\ \frac{{\bf i}\underaccent{\tilde}{\sigma}}{4}\partial_{p^{i}}\underaccent{\bar}{\bm b}_{\perp}\cdot\hat{\bm{\sigma}}\ +\ \frac{1}{4}({\bm{m}}\times\partial_{p^{i}}\underaccent{\bar}{\bm b}_{\perp})\cdot\hat{\bm{\sigma}}$, with $\underaccent{\bar}{\bm b}_{\perp}={\bm{m}}\times(\underaccent{\bar}{\bm b}_{\scriptscriptstyle SO}\times{\bm{m}})$. The intrinsic contribution to the SOT is then readily obtained from the trace of the first term on the RHS of Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems) with the spin operator, followed by momentum integration over the 2D Brillouin zone (BZ), and it reads
@@
${\bm{T}}^{\scriptscriptstyle(1)}_{int}\approx\frac{\hbar}{2}\int\displaylimits_{BZ}\!\!\frac{d^{2}\underaccent{\tilde}{p}}{(2\pi\hbar)^{2}}\left(f^{\scriptscriptstyle(eq)}_{\uparrow}-f^{\scriptscriptstyle(eq)}_{\downarrow}\right)(e{\bm{E}})\cdot\partial_{\bm{p}}\underaccent{\bar}{\bm b}_{\perp}.$ (12)
@@
Before considering the disorder correction, it is relevant to first recognize that ${\bm{E}}\cdot\partial_{\bm{p}}\underaccent{\bar}{b}_{\perp}^{i}=\partial_{\bm{p}}\cdot(\underaccent{\bar}{b}_{\perp}^{i}{\bm{E}})$, since under the additional assumption of the system to be degenerate, an application of the divergence theorem to the RHS of Eq.(12) then yields
@@
with $dl_{p}$ being the infinitesimal line element along the 1D Fermi surface for band $\sigma$ in the absence of SO interaction i.e., $FS_{\sigma}$. One shall add that if the Fermi surface intersects the BZ boundary, the line integrals over the portions of this boundary resulting from the application of the divergence theorem do not contribute to the RHS of Eq.(12). Indeed, since the 2D BZ is by definition a primitive unit cell of the reciprocal space over which $\hat{h}$ is periodic, such contributions will cancel out by pairs, from pairs of line segments on opposite sides of the BZ boundary, one being the translation of the other by a primitive reciprocal lattice vector. Hence, we have established that under the stated hypothesis, the intrinsic SOT, which according to Eq.(12) would be classified as a Fermi sea contribution, can equally be expressed as a Fermi surface one according to Eq.(13). This exactly parallels the classical result obtained by Haldane in the case of the non quantized part of the AHE conductivity [68], and is a further illustration that the decomposition of the linear response into Fermi sea and Fermi surface terms emphasized by the Kubo-Streda formula is neither unique nor particularly relevant.
Having established this non trivial result, we demonstrate now how our formalism delivers the disorder correction with striking efficiency. From the third term on the RHS of our Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems), we readily obtain the vertex correction to the interband Wigner function as
@@
Implications for the theory of orbital effects – Regarding the theory of out-of-equilibrium orbital angular momentum (OAM) in time reversal and inversion symmetric metals with weak SO interaction, of high current interest in connection with recent experiments [71, 69, 70], most Kubo formula based studies have focused on the intrinsic OHE conductivity [14, & references therein]. The intrinsic OHE current can be straightforwardly computed from the first term on the RHS of our Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems). Beyond the possible fragility of this current under disorder corrections [21], which can be very effectively assessed from the third term on the RHS of our Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems), and its debatable physical meaning as a non conserved quantity [95], we want to point out the importance of the gradient induced quantum coherence, as newly uncovered by the second term on the RHS of our Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems). It can be shown [96] that this is a source of OAM edge accumulation, only partially correlated to the OHE conductivity through the Berry connection, and which may be an important explanatory factor to the previously cited experimental observations.
@@
Conclusions – We provide a quantum kinetic theory of the linear response to an electric field slowly varying in time and space, for multiband electron systems in presence of weak scalar disorder. While our theory constitutes a formal and computational alternative to the Kubo formula in the stated context, with distinct advantages for finite systems and in going beyond the CBA, we have proven the equivalence of the two formulations at the ladder approximation level in the infinite and uniform case. The natural and unique decomposition of the linear response in our formalism, in term of an intraband response controlled by a transport equation in phase space, and interband coherences as local dependent quantities, provides new physical insights in the quantum structure of the linear response which are totally obfuscated in the Kubo formula. We have uncovered for the first time that spatial gradients in the intraband response on mesoscopic scales, which are ubiquitous in finite conductors, constitute a source of interband coherence. This opens a new research area as it suggests new ways to locally modulate current induced responses e.g., spin or orbital accumulation, by controlling the real space geometry or the boundary scattering properties. We have applied our approach to SOT theory for a simple model of 2DEG, recovering and generalizing with striking efficiency a major vertex correction, while uncovering a new instance of intrinsic response as a Fermi surface property. We also suggest possible implications for the theory of orbital effects, a subject of high current interest both from a fundamental and applied standpoint. Beyond analytical toy models, our formalism is equally applicable to realistic material models, up to ab-initio generated tight-binding Hamiltonian, through numerical resolution of the Boltzmann equation. We are already undertaking such studies in relation to SOT in van der Waals magnets with $C_{3}v$ symmetry [97], another subject of high current interest [18]. Finally, we shall stress that the simple disorder model considered in this Letter can be easily extended, for instance to include SO scattering or contributions from higher order diagrams beyond the ladder approximation, and that we also intend to address in a systematic manner the sub-leading contributions to the intraband response in a future work.